# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v3

@@ 5.3 Energy Efficiency Analysis
Figure 3: Energy consumption of DeSGDA and baselines on different datasets.
To assess the energy efficiency of DeSGDA, we use the metric from Zhu et al. (2022) and quantify the energy consumption for graph classification in the inference stage. Specifically, the graph domain adaption methods are evaluated on GPUs (NVIDIA A100), and the spiking-based methods are evaluated on neuromorphic chips (ROLLS Indiveri et al. (2015)) following Zhu et al. (2022). The results are shown in Figure 3, where we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs. Moreover, although the energy consumption of DeSGDA is slightly higher than DRSGNN due to additional computations required for domain adaptation, the performance improvement justifies the deployment of DeSGDA in low-power devices. Additionally, we present a comparison of training time and memory usage between DeSGDA and other graph domain adaptation methods. The results are detailed in Tables 9 and  10.
5.4 Ablation Study
We conduct ablation studies to examine the contributions of each component in the proposed DeSGDA: (1) DeSGDA w/o CA: It removes the adversarial distribution alignment module; (2) DeSGDA w/o PL: It removes the pseudo-label distilling module; (3) DeSGDA w/o CF: It removes the classification loss $\mathcal{L}_{S}$; (4) DeSGDA w/o TL: It utilizes the global thresholds on all nodes; (5) DeSGDA w/ PT: It deploys the adaptive perturbations Yin et al. (2022b) on source data for alignment; (6) DeSGDA w/ CL: It replaces the adversarial learning with the cross-domain contrastive learning Yin et al. (2023a).
Experimental results are shown in Table 2. From the table, we find that: (1) DeSGDA outperforms DeSGDA w/o CA, DeSGDA w/o PL, and DeSGDA w/o CF, demonstrating that the adversarial distribution alignment module effectively reduces domain discrepancies, ensuring well-aligned feature spaces between source and target domains. Additionally, the pseudo-label distillation module addresses variance in thresholds across domains, while the classification loss $\mathcal{L}_{S}$ enables DeSGDA to effectively learn from labeled source data and generalize to the target domain. (2) DeSGDA w/o TL shows lower performance compared to DeSGDA, showing that degree-aware thresholds, iteratively updated during training, resolve the issue of inflexible architecture in SGNs. By using these thresholds, DeSGDA effectively learns meaningful representations for nodes with various degrees. (3) DeSGDA outperforms DeSGDA w/ PT and DeSGDA w/ CL. We attribute this to the adaptive perturbations method (DeSGDA w/ PT) fail to help DeSGDA overcome domain discrepancies effectively. Additionally, the cross-domain contrastive learning method (DeSGDA w/ CL) is less effective at aligning the source and target distribution compared to adversarial learning. Furthermore, we provide ablation studies to examine the effect of directly replacing SGNs with commonly used Graph Neural Networks (GNNs) for generating representations for DeSGDA, and the results are shown in Tables 11, 12. More details about ablation results on other datasets are reported in Appendix H.3.
(a) Time Latency $\tau$
(b) Initial Threshold $V_{th}^{degree}$
(c) Balance Ratio $\alpha$
Figure 4: Hyperparameter sensitivity analysis of time latency $\tau$, initial threshold $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$ on PROTEINS.
@@ 5.5 Sensitivity Analysis
We conduct the sensitivity analysis of DeSGDA with respect to the impact of its hyperparameters: time latency $\tau$, degree threshold value $V_{th}^{degree}$ in SNNs, and balance ratio $\alpha$, which plays a crucial role in the performance of DeSGDA. In particular, $\tau$ controls the number of SNNs propagation steps; $V_{th}^{degree}$ determines when a neuron fires; $\alpha$ governs the changing ratio of degree-aware thresholds.
Figure 4 illustrates how $\tau$, $V_{th}^{degree}$, and $\alpha$ affects the performance of DeSGDA on the PROTEINS dataset. More results on other datasets are shown in Appendix H.4. We vary $\tau$ within the range of $\{5,6,7,8,9,10\}$, $V_{th}^{degree}$ in $\{0.05,0.1,0.2,0.5,1.0,2.0,5.0\}$, and $\alpha$ in $\{0.1,0.3,0.5,0.7,0.9\}$. From the results, we observe that: (1) The performance of DeSGDA in Figure 4(a) generally exhibits an increasing trend at the beginning and then stabilizes when $\tau$ is greater than 8. We attribute this to smaller values of $\tau$ potentially losing important information for representation, while larger values significantly increase model complexity. To balance performance and complexity, we set $\tau$ to 8 as default. (2) Figure 4(b) indicates an initial increase followed by a decreasing trend in performance as $V_{th}^{degree}$ increases. This trend occurs because a lower threshold may trigger more spikes for high-degree nodes, leading to a drastic change in threshold, which degrades performance. Conversely, a higher threshold for low-degree nodes could result in fewer spikes, affecting model’s ability to process information. Thus, we set $V_{th}^{degree}$ to 0.5 as default. (3) From Figure 4(c), we find that the performance of DeSGDA initially increases and then decreases as $\alpha$ increases. The potential reason is that smaller $\alpha$ may delay the updating of the threshold, leading to performance degradation. Contrarily, a larger $\alpha$ tends to introduce more spikes that change dynamically at each step, resulting in instability in the model’s performance. Therefore, we set $\alpha$ to 0.5 as default.
@@ 6 Conclusion
In this paper, we propose the problem of spiking graph domain adaptation and introduce a novel framework, DeSGDA, for graph classification. This framework enhances the adaptability and performance of SGNs through three key aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation. Our approach enables expressive information capture through degree-dependent spiking thresholds, aligns feature distributions via adversarial training, and utilizes pseudo-labels to leverage unlabeled data effectively. Extensive experimental validation across benchmark datasets demonstrates that DeSGDA surpasses existing methods in accuracy while maintaining efficient energy consumption, making it a promising solution for advancing the domain adaptation capabilities of spiking graph networks. In the future, we will apply SGNs in source-free domain adaptation and domain generalization scenarios.
Appendix A Proof of Hypothesis 1
@@
$h_{i}=h_{i}+\sum_{j\in N(i)}w_{ij}h_{j}.$ (9)
Therefore, we have the expectation:
@@
$\mathbb{E}(h_{i})\sim\mathcal{N}\left((1+\sum_{j\in N(i)}w_{ij})\mu,(1+\sum_{j\in N(i)}w_{ij})\sigma^{2}\right).$ (11)
From the results, we observe that node $i$ follows a normal distribution with a mean of $(1+\sum_{j\in N(i)}w_{ij})\mu$, determined by the aggregated weights of its neighboring nodes. To provide a more intuitive understanding, we visualize the aggregated neighbor weights of GCN Kipf and Welling [2017] and GIN Xu et al. [2018] in Figure 5. The results show that as the degree increases, the aggregated weights also increase progressively. Consequently, high-degree nodes tend to follow a normal distribution with a higher mean and variance. In other words, nodes with higher degrees accumulate greater signals, making them more likely to trigger spiking. Based on this, we propose assigning higher thresholds to high-degree nodes and lower thresholds to low-degree nodes.
Another observation is that methods that normalize neighbor weights to 1 (e.g., GAT Veličković et al. [2017], GraphSAGE Hamilton et al. [2017]) still result in aggregated features following the same normal distribution. This normalization diminishes the ability to distinguish between nodes with varying degrees, ultimately degrading performance. This explains why, when using GAT as the backbone of DeSGDA, the performance is the weakest.
Figure 5: Visualization of degree-aware thresholds and aggregation weights on PROTEINS and NCI1.
Appendix B Proof of Theorem 2
@@
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right),$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
@@
where $(a)$ results from Shen et al. [2018] Theorem 1 with the assumption $\max(||h||_{Lip},\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})})\leq C_{f}C_{g}$, $D\in\{S,T\}$. Similarly, we obtain:
@@
$+\min\left(|\epsilon_{S}(h,\hat{h}_{S})-\epsilon_{S}(h,\hat{h}_{T})|,|\epsilon_{T}(h,\hat{h}_{S})-\epsilon_{T}(h,\hat{h}_{T})|\right)$
The $\hat{h}$ above is the abbreviation of $\hat{h}(x)$, which means the input is the continuous feature. Provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}$, and motivated by Yin et al. [2024a], we have:
@@
Furthermore, motivated by the techniques given by Bartlett and Mendelson [2002], we have:
@@
From McDiarmid’s inequality McDiarmid et al. [1989], with probability at least $1-\delta$, we have:
@@
where $\epsilon_{i}$ is the Rademacher variable. Combining Eq. 23 24 26, we have:
@@
Appendix C Proof of Theorem 3
Theorem 3 Under the assumption of Theorem 1, we further assume that there exists a small amount of i.i.d. samples with pseudo labels $\{(G_{n},Y_{n})\}_{n=1}^{N^{\prime}_{T}}$ from the target distribution $\mathbb{P}_{T}(G,Y)$ $(N^{\prime}_{T}\ll N_{S})$ and bring in the conditional shift assumption that domains have different labeling function $\hat{h}_{S}\neq\hat{h}_{T}$ and $\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})}=C_{h}\leq C_{f}C_{g}(D\in\{S,T\})$ for some constant $C_{h}$ and distance measure $\eta$, and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality holds:
$+2C_{f}C_{g}W_{1}\left(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G)\right)+\omega^{\prime}.$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega^{\prime}=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:

### v4

@@ 5.4 Energy Efficiency Analysis
To assess the energy efficiency of DeSGraDA, we use the metric from (Zhu et al., 2022) and quantify the energy consumption for graph classification in the inference stage. Specifically, the graph domain adaptation methods are evaluated on GPUs (NVIDIA A100), and the spiking-based methods are evaluated on neuromorphic chips (ROLLS (Indiveri et al., 2015)) following (Zhu et al., 2022). The results are shown in Figure 4(a), where we find that compared with traditional graph domain adaptation methods, the spike-based methods (DeSGraDA and DRSGNN) have significantly lower energy consumption, demonstrating the superior energy efficiency of SGNs. Moreover, although the energy consumption of DeSGraDA is slightly higher than DRSGNN due to additional computations required for domain adaptation, the performance improvement justifies deploying DeSGraDA in low-power devices. Additionally, we present a comparison of training time and memory usage between DeSGraDA and other GDA methods, and the results are detailed in Tables 5 and  6.
(a) Energy Consumption
(b) Latency Step $T$
(c) Initial Threshold $V_{th}^{degree}$
Figure 4: (a) Energy efficiency analysis and (b), (c) hyperparameter sensitivity analysis of latency step $T$ and initial threshold $V_{th}^{degree}$ on the PROTEINS dataset.
@@ 5.5 Sensitivity Analysis
We conduct the sensitivity analysis of DeSGraDA to investigate the impact of key hyperparameters: latency step $T$ and degree threshold $V_{th}^{degree}$ in SGNs. Specifically, $T$ controls the number of SGNs propagation steps, and $V_{th}^{degree}$ determines the firing threshold of each neuron based on node degree.
Figure 4(b) and 4(c) illustrates how $T$ and $V_{th}^{degree}$ affects the performance of DeSGraDA on the PROTEINS dataset. More results on other datasets are shown in Appendix K.4. We vary $T$ in $\{5,6,7,8,9,10\}$ and $V_{th}^{degree}$ in $\{0.05,0.1,0.2,0.5,1.0,2.0,5.0\}$. From the results, we observe that: (1) The performance of DeSGraDA in Figure 4(b) generally exhibits an increasing trend at the beginning and then stabilizes when $T>9$. We attribute this to smaller values of $T$ potentially losing important information for representation, while larger values significantly increase model complexity. To balance effectiveness and efficiency, we set $T=9$ as default. (2) Figure 4(c) indicates an initial increase followed by a decreasing trend in performance as $V_{th}^{degree}$ increases. This trend arises because a lower threshold may cause excessive spiking for high-degree nodes, leading to unstable representation, while a higher threshold may suppress spiking for low-degree nodes, reducing information flow. Accordingly, we set $V_{th}^{degree}$ to 0.2 as default.
@@ 6 Conclusion
In this paper, we propose the problem of spiking graph domain adaptation and introduce a novel framework DeSGraDA for graph classification. DeSGraDA enhances the adaptability and performance of SGNs through three key aspects: degree-conscious spiking representation, temporal distribution alignment, and pseudo-label distillation. DeSGraDA captures expressive information via degree-dependent spiking thresholds, aligns feature distributions through temporal dynamics, and effectively exploits unlabeled target data via pseudo-label refinement. Extensive experiments on benchmark datasets demonstrate that DeSGraDA surpasses existing methods in accuracy while maintaining energy efficiency, showcasing its potential as a strong solution for DA in SGNs. In the future, we plan to extend DeSGraDA to source-free and domain-generalization scenarios.
Appendices
Appendix A DA Bound for Graphs
DA Bound for Graphs. Due to the DA theory is agnostic to data structures and encoders, You et al. [2023] directly rewrite it for graph-structured data ($G$) accompanied with graph feature extractors ($f$) as follows, and the covariate shift assumption is reframed as $\mathbb{P}_{S}(Y|G)=\mathbb{P}_{T}(Y|G)$.
Theorem 3
[You et al., 2023] Assume that the learned discriminator is $C_{g}$-Lipschitz continuous as in [Redko et al., 2017], and the graph feature extractor $f$ is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, with probability at least $1-\delta$ the following inequality holds:
$\epsilon_{T}(h,\hat{h})\leq\ \hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log\left(\frac{eN_{S}}{d}\right)+\frac{1}{N_{S}}\log\left(\frac{1}{\delta}\right)}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
where $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$ is the (empirical) target risk, $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ is the (empirical) source risk, $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$, and $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$ is the first Wasserstein distance [Villani et al., 2008].
Appendix B Spiking Neural Networks
Spiking Neural Networks (SNNs) are brain-inspired models that communicate through discrete spike events rather than continuous-valued activations Maass [1997], Gerstner and Kistler [2002], Bohte et al. [2000]. This design provides significant advantages in temporal information processing and energy efficiency. Different from traditional neural networks, SNNs emulate biological mechanisms such as membrane potential integration, threshold-triggered firing, and post-spike resetting Cao et al. [2015], Zhu et al. [2022]. To capture these biological dynamics, SNNs employ neuron models that mathematically describe the temporal evolution of membrane potentials and the conditions for spike generation Lagani et al. [2023]. A widely used neuron model in SNNs is the Leaky Integrate-and-Fire (LIF) model Tal and Schwartz [1997], Lansky and Ditlevsen [2008], which operates through three fundamental stages:
(1) Integrate: At each lantency step $t$, the membrane potential $V[t]$ is updated by integrating the input current $I[t]$ and applying a decay to the previous potential $V[t-1]$:
$V[t]=\lambda V[t-1](1-S[t-1])+I[t]$ (10)
where $\lambda\in(0,1)$ is the decay factor that controls the leakage rate, and $S[t-1]$ is the binary spike indicator from the previous time step.
(2) Fire: A spike is emitted when the membrane potential exceeds the threshold $V_{th}$:
$S[t]=H(V[t]-V_{th}),$ (11)
where $H(\cdot)$ denotes the Heaviside step function, a non-differentiable function defined as:
$H(x)=\begin{cases}1,&x\geq 0,\\ 0,&\text{ otherwise. }\end{cases}$ (12)
(3) Reset: After a spike is emitted, the membrane potential is reset according to the following rule:
$V[t]=(1-S[t])V[t]+S[t]V_{reset},$ (13)
where $V_{reset}$ denotes the resting potential, typically set to zero.
Due to the non-differentiability of $H(\cdot)$, surrogate gradient methods are commonly employed to approximate its derivative during backpropagation, enabling gradient-based optimization in deep SNN architectures Bohte et al. [2000], Sun et al. [2024].
Appendix C Proof of Proposition 1
@@
$h_{i}=h_{i}+\sum_{j\in N(i)}w_{ij}h_{j},$ (14)
where $N(i)$ is the neighbor set of node $i$. Therefore, we have the expectation:
@@
$\mathbb{E}(h_{i})\sim\mathcal{N}\left((1+\sum_{j\in N(i)}w_{ij})\mu,(1+\sum_{j\in N(i)}w_{ij})^{2}\sigma^{2}\right).$ (16)
From the results, we observe that node $i$ follows a normal distribution with a mean of $(1+\sum_{j\in N(i)}w_{ij})^{2}\mu$, determined by the aggregated weights of its neighboring nodes. To provide a more intuitive understanding, we visualize the aggregated neighbor weights of GCN [Kipf and Welling, 2017] and GIN [Xu et al., 2018] in Figure 2. The results show that as the degree increases, the aggregated weights also increase progressively. Consequently, high-degree nodes tend to follow a normal distribution with a higher mean and variance. In other words, nodes with higher degrees accumulate greater signals, making them more likely to trigger spiking. Based on this, we propose assigning higher thresholds to high-degree nodes and lower thresholds to low-degree nodes.
Appendix D Proof of Theorem 2
@@
[…]
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h}(\mathbf{S}))=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(\mathbf{S}_{n})-\hat{h}(\mathbf{S}_{n})|$ and $\epsilon_{T}(h,\hat{h}(\mathbf{X}))=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X}))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X}))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X}))|\right)$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is the probability matrix with:
@@
where $(a)$ results from [Shen et al., 2018] Theorem 3 with the assumption $\max(||h||_{Lip},\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})})\leq C_{f}C_{g}$, $D\in\{S,T\}$. Similarly, we obtain:
@@
The $\hat{h}$ above is the abbreviation of $\hat{h}(x)$, which means the input is the continuous feature. Provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}$, and motivated by [Yin et al., 2024], we have:
@@
Furthermore, motivated by the techniques given by [Bartlett and Mendelson, 2002], we have:
@@
From McDiarmid’s inequality [McDiarmid et al., 1989], with probability at least $1-\delta$, we have:
@@
where $\epsilon_{i}$ is the Rademacher variable. Combining Eq. 28 29 31, we have:
@@
Appendix E Generalization Bound with Pseudo-label Distillation Module
Theorem 4
Under the assumption of Theorem 3, we further assume that there exists a small amount of i.i.d. samples with pseudo labels $\{(G_{n},Y_{n})\}_{n=1}^{N^{\prime}_{T}}$ from the target distribution $\mathbb{P}_{T}(G,Y)$ $(N^{\prime}_{T}\ll N_{S})$ and bring in the conditional shift assumption that domains have different labeling function $\hat{h}_{S}\neq\hat{h}_{T}$ and $\max_{G_{1},G_{2}}\frac{|\hat{h}_{D}(G_{1})-\hat{h}_{D}(G_{2})|}{\eta(G_{1},G_{2})}=C_{h}\leq C_{f}C_{g}(D\in\{S,T\})$ for some constant $C_{h}$ and distance measure $\eta$, and the loss function bounded by $C>0$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$, and provided the spike training data set $S_{n}=\{(\mathbf{X}_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\}_{i\in[n]}$ drawn from $\mathcal{D}^{s}$, with probability at least $1-\delta$ the following inequality holds:
[…]
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min\left(|\epsilon_{S}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{S}(h,\hat{h}_{T}(\mathbf{X})))|,|\epsilon_{T}(h,\hat{h}_{S}(\mathbf{X})))-\epsilon_{T}(h,\hat{h}_{T}(\mathbf{X})))|\right)$, $\epsilon_{i}$ is the Rademacher variable and $p_{i}$ is the $i^{th}$ row of $\mathbf{P}$, which is defined in Theorem 2.